A rectangular rug has a perimeter of 140 feet. Its length is 10 feet longer then its width. What are the demenisions?

Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

2L+2W=140
2(W+10)+2W=140
2W+20+2W=140
4W=140-20
4W=120
W=120/4
W=30 FEET FOR THE WIDTH.
L=30+10=40 FEET FOR THE LENGTH.
PROOF
2*30+2*40=140
60+80=140
140=140

Answer by tutorcecilia(2152)

Perimeter = 2(length+width)
Let length = width +10
Let width = w
.
140=2[(w+10)+(w)]
140/2=2w+10
70=2w+10
70-10=2w
60=2w
60/2=w
30=w
.
Width = 30
Length=w+10=30+10=40
.
Therefore
140=2(40+30)
140=2(70)
140=140 [checks out]